Week of 060410

Open Actions from last week:

Chair: James

On Call: James/Maarten

Monday:

Log:
  • FTS006 does not have correct root password for sysadmins
  • LFC_READDIR errors - no procedures

New Actions:

  • check state of fts006 (james) - DONE
  • write op/ssyadmin procedures for gridlfc cluster -DONE
  • Take out FNAL and monitor traffic response - DONE

Discussion:

  • current running at 1.3Gb - still many sites need to reach their throughput. Also, nned to investigate if we have a bottleneck at 1.4Gb at CERN.

Tuesday:

Log: DNS outage for 1.5hrs. Caused all traffic to stop and affected other services (SPMA, etc...). second (small) dip at ~21:00 - unexplained

New Actions:

Discussion:

  • No announcement was made (and no CS rep present at morning meeting) of intervention
  • Taking out FNAL made no difference - seems not to be a bottleneck due to #transfers at 1.4G

Wednesday

Log: Slow fade of transfers at 9pm. Suspected link to cleanup of FTS DB

Actions:

  • LFC move to new hardware of lfc-dteam-test,lfc-cms-test,lfc-atlas-test - James -DONE
Discussion:

Thursday

Log: Reached 1.6GB after DB cleanup

Actions:

  • Check with sophie status of message re: RLS shutdown

Discussion:

Friday

Log:

Actions:

Discussion:

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2006-04-13 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback